40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Aug 10 11:47:19 2017, awade, DailyProgress, BEAT, Improving pre-cavity BN detector 20170810_PreCavBNDetectorConfig.pdf
    Reply  Fri Aug 11 16:58:42 2017, awade, koji, DailyProgress, BEAT, Improving pre-cavity BN detector 
Message ID: 1854     Entry time: Fri Aug 11 16:58:42 2017     In reply to: 1852
Author: awade, koji 
Type: DailyProgress 
Category: BEAT 
Subject: Improving pre-cavity BN detector 

Koji came down to the lab to have a look at the pre-cavity BN dector.  With a might tigher lens f = 37 mm he was able to get a DC voltage of - 2.59 V for an total input power of 577 µW.  We should expect about 4.00 V for this power, but this is what we get.  We couldn't see a beat note until Koji started changing the laser crystal temperature on the south laser controller.  The set point temperature was changed from 34.4961 C to 47.0749 C. Slow laser voltage offset channels were set to 0.6901 V for the south path and 3.8294 V for the north path.  With these settings we were able to see a BN at 216 MHz of -50 dBm. This is sufficient for the diagnostic purpose of this detector.

For reference, here is a configuration summary of the lasers and controls at present:

BN = -50 dBm @ 216 MHz

South laser configuration
  Setting
ADJ 0
DC 2.04 A
DPM 0.00V
NE ON
LD ON
Display 5
DT 28.8 C
DTEC +0.6V
LT 47.2 C
LTEC -0.1 V
T +47.0749
Pwr 160 (not calibrated)

South slow frequency voltage offset = 0.6901 V

 

North laser configuration
  Setting
ADJ 0
DC 2.08 A
DPM -0.00V
NE ON
LD ON
Display 1
DT 22.3 C
DTEC +0.6V
LT 40.2C
LTEC 0.0 V
T +26.4650 C
Pwr 68 (not calibrated)

South slow frequency voltage offset = 3.8294 V

Todo: make a BN laser north vs laser south map similar to what koji did in this this post 40m:3759.  This should make the good operating parameter regiem obvious in the future.

ELOG V3.1.3-